# Ultracold Interactions between Ions and Polar Molecules

### v1

@@ Abstract
We discuss a platform for observing and controlling the interactions between atomic ions and a quantum gas of polar molecules in the ultracold regime. This approach is based on the combination of several recently developed methods in two so-far complementary research domains: ion-atom collisions and studies of ultracold polar molecules. In contrast to collisions between ions and ground-state atoms, which are dominated by losses due to three-body recombination (TBR) already at densities far below those typical for quantum degenerate ensembles, our proposal makes use of polar molecules, their rich level structure, and sensitivity to electric fields to design effective interaction potentials where ion-neutral TBR losses and molecule-molecule losses due to sticky collisions are strongly suppressed. This may enable access to the deep quantum regime of interaction with a broad range of applications including the potential formation of novel many-body self-bound states, quantum simulations, and quantum chemistry between polyatomic molecules.
ion-atom interactions, molecular quantum gases, quantum chemistry, many-body physics
[…]
1 Introduction
@@
Controlling collisions between ions and neutral atoms in a regime where quantum effects are dominant is a long-standing goal in the field of ion-atom interactions [1, 2, 3]. This comparatively new and interdisciplinary research domain holds great promise in advancing or enabling a broad spectrum of applications in areas including quantum many-body and polaron physics [4, 5, 6, 7], charge mobility [8, 9], quantum simulations (QS) [10, 11], quantum chemistry (QC) [12, 13, 14, 15, 16, 17], and quantum information processing (QIP) [18, 19, 20]. Most of these applications either strongly benefit from or require ultralow temperatures where collisional processes are dominated by the contribution of the lowest partial wave, the so-called s-wave regime. During the last years, enormous progress has been made towards this goal. In particular, recent advances have enabled several key achievements such as the elimination of micromotion-induced heating in generic compound optical traps [21], sympathetic cooling below the Doppler limit and close to the s-wave regime [22, 17, 21], and the controlled association of ion-atom Feshbach molecules [17].
@@
However, some substantial challenges remain an obstacle in view of many envisioned applications which require access to the deep quantum regime characterized by a combination of quantum degenerate atomic gases and ions in or close to the motional ground state. This is largely due to inelastic or reactive collisions which dominate already at moderate neutral densities [23, 21, 24] even for spin-polarized fermionic ensembles [22, 17] and lead to loss of the ion. They originate from the comparatively large characteristic range of the interaction potential of typically $R^{\star}\sim 100~{}\textrm{nm}$ between an electric charge and the induced dipole moment of the atoms exposed to the electric field of the ion. For ground-state atoms located at the distance $R$ from the ion, this attractive interaction potential scales as $1/R^{4}$ and manifests itself in the form of strong enhancement of ion-atom-atom three-body recombination recombination processes (TBR) [23, 25, 26, 27]. At the same time, the attractive interaction is also responsible for a displacement of the ion from the node of the radiofrequency (rf) quadrupole field used to generate the confining potential for the ion, facilitating energy transfer from the rf field to the center of mass motion of the ion-atom system whenever they are allowed to undergo collisions [28].
@@
This so-called micromotion-induced heating has been identified as one of the principal obstacles to reaching the s-wave regime in conventional hybrid traps utilizing rf fields to confine the ions [28]. In order to circumvent these limitations, several strategies have been brought forward making use of different systems and mechanisms such as Rydberg atom-atom interactions [29, 30], Rydberg dressing [31], restricting the studies to few specific combinations with an extremely low reduced mass [28, 22, 17], and bichromatic traps for confining ion and atoms [21, 32]. While all these strategies have enabled recent breakthroughs, the combination of ions with quantum degenerate atomic ensembles which are characterized by large phase-space densities on the order of 1 and comparatively high densities of typically $n\sim 10^{15}~{}\textrm{cm}^{-3}$, is still out of reach. This is a consequence of the strong enhancement of TBR in vicinity of an ion in conjunction with the scaling of the related loss rates as $n^{2}$ resulting in drastically reduced ion lifetimes expected for the required densities [23, 21, 17].
@@
In parallel to the rapid evolution of studies focusing on ion-atom combinations, the complementary field of molecular quantum gases has seen a similarly impressive development. While the focus of the field shifted from chemically reactive bialkali mixtures investigated in first ground-breaking experiments [33, 34] to nonreactive combinations and directly laser cooled molecular species, the surprisingly limited lifetimes observed for all species emerged as one of the main challenges. For non-reactive combinations, the underlying losses have prevented the implementation of evaporative cooling, and until recently, hampered the creation of quantum degenerate gases. The exact mechanism behind these losses remains unclear and is a subject of ongoing debates [35, 36], but it is likely related to so-called sticky collisions, which occur at short intermolecular distances and involve the formation of metastable collisional complexes [37, 38, 35, 36]. With the development of methods to prevent the molecules from reaching the short range [39, 40, 41, 42, 43], this obstacle has recently been overcome [44, 45]. Such shielding methods rely on engineered potential barriers created through interactions with microwave, optical or static electric fields. They have enabled efficient evaporative cooling [46, 47, 48], and the first observations of molecular Fermi gases [49, 46] and Bose-Einstein Condensates (BECs) [48] have now become a reality.
@@
Here, we discuss the possibility to combine the methodology and advantageous properties of the respective systems employed in both fields in order to extend experimental studies of ion-molecule interactions at ultralow temperatures towards quantum gases of polar molecules. First, we give a brief overview of the recent developments in the fields of ultracold molecules and ion-atom interactions, discuss how existing techniques can be used to realize a suitable hybrid platform, and derive the expected properties and behavior of polar molecules exposed to the electric field of a stationary ion. In the subsequent sections, we investigate how the gain of additional degrees of freedom and the resulting interplay of attractive and repulsive long-range and short-range interactions can be used to engineer mechanisms for suppressing undesired processes, both in dilute and in dense ensembles of polar molecules. Lastly, we discuss potential applications of the proposed hybrid ion-molecule system in studies of quantum-chemical reactions, many-body phenomena, and dynamical processes in molecular quantum gases as well as their control.
2 Experimental toolbox for hybrid trapping of ions and neutral molecules
In parallel to these developments in the field of ultracold molecular gases, first experiments for studying ultracold ion-atom collisions have been realized [12, 13]. Their aim was to combine vastly different methodologies for preparing and controlling atoms and ions at ultralow temperatures in a common two-species apparatus. Such hybrid setups using radiofrequency (rf)-driven Paul traps to confine ions and (magneto-) optical traps for preparing and controlling ultracold atomic ensembles have proven to be a versatile platform and have been successfully demonstrated for many ion-atom combinations [12, 13, 50, 14, 51, 52, 16, 24, 53, 54, 55, 56, 17, 22]. More recently, it was shown that this concept can be extended to combinations of several species of atomic ions and multi-species atomic magneto-optical traps which can be used to study ion-atom collisions at low temperatures and in principle support the association of bialkali molecules [56]. Among the many important developments in this field, we point out two specific examples that are most relevant in view of the proposed extension of the study of ion-neutral interactions to ensembles of molecular quantum gases, that is, highly optimized rf-based hybrid traps and all-optical ion-atom traps. In the former case, for favorable ion-atom combinations featuring low reduced massed and long-range induction coefficients, optimized Paul traps in conjunction with advanced schemes for the detection and compensation of stray electric fields have allowed to reduce the collision energy to a range where non-classical effects become visible [22, 17]. In the latter case, similarly optimized setups were also combined with bichromatic optical traps capable of simultaneously confining ions and atoms [21, 32]. In particular, using optical traps provides the advantage that micromotion-induced heating can be completely eliminated and that this method is applicable to generic combinations of species provided that an effective optical trapping potential can be engineered [57].
Figure 1: (a) Schematic of an ion-molecule hybrid trap combining a magnetic trap, optical dipole traps, and an integrated trap for ions for investigating and controlling the interactions between single ions and neutral molecular quantum gases. (b) An ion immersed into the neutral cloud introduces a source of static electric fields, aligning the permanent molecular electric dipoles, depending on the distance $R$. This provides access to a new system with a rich hierarchy of competing tunable interactions spanning 1/$R^{6}$ (induced dipole – dipole, attractive), 1/$R^{4}$ (charge – induced dipole, attractive or repulsive: depending on the quantum numbers $J,m_{J}$), 1/$R^{3}$ (dipole – dipole, attractive for side-by-side or repulsive for head-to-tail collisions), 1/$R^{2}$ (charge – dipole, attractive or repulsive: determined by the electric field and $J,m_{J}$). The transition from the 1/$R^{4}$ to the 1/$R^{2}$ potential occurs above the characteristic electric field $\mathcal{E}^{\ast}$.
In principle, this is the case for any combination of ions and neutral atoms or molecules possessing a non-zero dipole moment.
In addition to ensembles of bialkali molecules such as those given in Table 1, the same concept can be applied to experiments on the level of individual particles, e.g. single molecules in optical tweezers [58, 59, 60, 61] and to other classes of polar molecules, e.g. directly laser-coolable CaF [62, 63, 61] and BaF [64], NaCs [58], RbCs [59], and polyatomic molecules such as CaOH [60] or CaOCH3 [65]. For clarity, we will focus our discussion on the exemplary ensembles of ground-state NaK and Ba+ ions for the following reasons. Firstly, molecular quantum (-degenerate) gases with a high phase-space density have been recently achieved for NaK [66, 46], and secondly, both the ions and the molecules can be routinely confined and prepared in standard optical dipole traps [67, 68, 69, 70, 66, 46]. Furthermore, as we will discuss in the remainder of this section, NaK molecules prepared in the rovibronic ${}^{1}\Sigma^{+}$ ground state have very similar properties to common alkali species regarding mass and dynamic dipole polarizability $\alpha_{d}(\omega)$ with $\omega$ denoting the frequency of the interacting optical field. This allows for detailed comparisons of the expected behavior of the proposed ion-molecule system with Ba+-Rb mixtures which have been extensively studied both in presence of micromotion-induced heating encountered in rf-hybrid [23, 25, 24] and in fr-free optical traps where micromotion is eliminated [21]. Lastly, near-infrared (NIR) far-off resonant optical traps operated at a wavelength of 1064 nm, have been successfully used both to trap NaK and to demonstrate a generic approach to studying ion-atom collisions, first realized for Ba+-Rb mixtures [21]. To this end, bichromatic optical potentials are created by overlapping dipole traps of different wavelengths.
For a direct comparison with previous experiments we use the parameters employed in [21], that is, optical dipole traps with wavelengths of 532 nm (VIS ODT) and 1064 nm (NIR ODT), optical powers $P_{VIS}=0.47~{}\textrm{W}$ and $P_{NIR}=0.13~{}\textrm{W}$ (adjusted for typical temperatures of NaK quantum gases), matched $1/e^{2}$ waist radii $w_{0}^{VIS}=w_{0}^{NIR}=3.8\times 10^{-6}~{}\textrm{m}$, a residual stray electric field of $\mathcal{E}_{str}=10~{}\textrm{mV}~{}\textrm{m}^{-1}$, and the axial secular frequency of the linear Paul trap of $\omega_{x}/(2\pi)=12~{}\textrm{kHz}$, respectively. We then calculate the bichromatic potential suitable for overlapping a barium ion laser cooled close to the Doppler temperature $T_{D}^{Ba}=365~{}\mu\textrm{K}$ with an ensemble of ground-state NaK molecules prepared at $T^{NaK}=400~{}\textrm{nK}$. The NaK molecules experience a repulsive potential from the VIS ODT due to its blue detuning with respect to the energetically lowest dipole-allowed optical transition $X^{1}\Sigma^{+}\leftrightarrow A^{1}\Sigma^{+}$. To maintain the initial trap depth $U_{0}^{NaK}/(k_{B})\approx 4~{}\mu\textrm{K}$, the intensity of the NIR ODT has to be increased compared to its initial value in order to achieve the desired differential light shift. Despite the substantially more complex energy level structure of molecules in comparison with alkali atoms, the resulting potentials calculated in the far-off-resonant limit [71] and using the dynamic dipole polarizability of NaK [72], that are shown in Fig.2 are quantitatively very similar to those obtained for the previously demonstrated case of Ba+-Rb in terms of trap depths and ODT parameters. This suggests that the protocols for overlapping ions and atoms demonstrated in [21] as well as proposed methods for further improving the trapping performance and realizing more complex trapping configurations [69, 57, 70] are readily applicable to molecules. The expected compatibility to such an extent may seem surprising at first glance because, in general, neutral atoms and polar molecules may have vastly different polarizabilities. This is indeed the case for the static polarizability $\alpha_{0}$ which is a pivotal point of the proposed experiments and is discussed in detail in section 3. However, in the optical and near-infrared frequency domains where the detunings with respect to rotational transitions are several orders of magnitude larger than those for optical transitions, all $\alpha_{d}(\omega)$ for the bialkali molecules in the rovibronic ground state as listed in Table 1 vary by less than a factor of 2 and are of the same order of magnitude as those of common alkali atoms [72]. It is this robustness of the dynamic polarizability that facilitates the creation of molecules from individually prepared atomic species which are typically all confined in the same far-detuned optical dipole traps. At the same time, even within the same electronic and vibrational ground state manifold, i.e. $\left|\textit{X}^{1}\Sigma^{+},\textrm{v}=0\right\rangle$, $\alpha_{d}(\omega)$ exhibits a substantial dependence on the rotational quantum numbers $J$ and $m_{J}$ [73].
Figure 2: Effective bichromatic radial (orthogonal to the ODT’s wavevector) potentials (black solid lines) for Ba+ (in the electronic ground state $6^{2}S_{1/2}$) and NaK (in the rovibronic ground state $\left|\textit{X}^{1}\Sigma^{+},\textrm{v}=0,\textrm{J}=0\right\rangle$) exposed to a composite optical dipole trap superimposing ODTs at 1064 nm (NIR, red dashed-dotted line) and at 532 nm (VIS, green dashed line). The effective potential of NaK is enlarged by a factor of 100 for visibility. The Paul and optical dipole trap parameters are adopted from experiments with Ba+-Rb mixtures, accounting for the properties of NaK at a temperature of $T^{NaK}=400~{}\textrm{nK}$.
Exploiting the similarities between molecules and alkali atoms interacting with far-detuned optical fields, a prototypical experimental setup could be realized as shown in Fig.1. Its central features include (i) a Paul trap optimized for ion-neutral experiments, that is, optical access through high numerical aperture objectives beneficial for stray electric field compensation and focusing of the dipole trap beam(s) and the correspondingly designed ion-electrode spacing $(\sim 1~{}\textrm{mm})$, (ii) magnetic field coils required for the magneto-association of Feshbach molecules, as well as (iii) a bichromatic optical dipole trap. Similar Paul traps have been demonstrated and feature suppression of stray electric fields down to $\sim 10^{-3}~{}\textrm{V m}^{-1}$ [21, 17, 74, 70], very low anomalous heating rates [68], and typical trap depths on the order of $k_{B}\times 10^{4}~{}\textrm{K}$, which is sufficiently deep to recapture ionic products.
The above considerations suggest that the proposed hybrid platform is suitable for investigations of ultracold ion-molecule interactions, such as precise studies of controlled quantum chemistry which are discussed in section 4. In the following sections, we focus on the expected behavior of the system and discuss several potential applications of the proposed hybrid platform.
Table 1: Characteristic properties of the charge-induced dipole long-range interaction between common heteronuclear bialkali molecules prepared in the rovibronic ground state and a 138Ba+ ion. The static dipole polarizabilities $\alpha_{0}$ and permanent electric dipole moments (PDM) are taken from [72]. For comparison, 87Rb and 6Li are given as a reference for ion-atom systems, representing a typical and one of the most favorable combinations with respect to reaching the s-wave scattering limit, respectively.
  • PDM $\alpha_{0}$ $R^{\star}$ $E^{\star}/k_{\textrm{B}}$
    (D) $(10^{-37}~{}\textrm{C m}^{2}\textrm{V}^{-1})$ $(10^{-6} \textrm{m})$ (nK)
    LiCs 5.59 312 25.9 $0$.00521
    NaCs 4.69 709 40.1 $0$.00206
    LiRb 4.18 152 16.2 $0$.0165
    LiK 3.58 093.6 10.0 $0$.0703
    NaRb 3.31 294 23.6 $0$.00711
    NaK 2.78 152 14.2 $0$.0281
    KCs 1.84 208 22.2 $0$.00644
    RbCs 1.25 177 21.6 $0$.00614
    KRb 0.62 018.80 06.21 $0$.0952
    LiNa 0.57 001.64 01.12 0007.86
    Rb 0.00 000.0526 00.295 0052.3
    Li 0.00 000.0270 00.0694 8760
3 Polar molecules in the electric field of an atomic ion
@@
Figure 3: (top panel) Magnitude of the electric field emanating from a single ion as a function of its distance to an approaching neutral molecule. The shaded area indicates the distance range of Fig.5, where charge exchange reactions are exothermic. The corresponding electric potential experienced by the molecules becomes extremely large ($>100~{}\textrm{kV cm}^{-1}$) resulting in almost complete alignment of the molecular dipole moment ($d(\mathcal{E})>0.9~{}d_{0}$). The grey dashed line denotes the characteristic electric field of NaK, $\mathcal{E}^{\star}$(NaK)$=d_{0}/B_{0}=2.04~{}\textrm{kV cm}^{-1}$. (bottom panel) Numerically calculated energy levels of the NaK molecule for the lowest rotational states $\left|J=0,~{}1,~{}2,~{}3;m_{J}\right\rangle$ (considering a maximal $J_{max}=10$) within the vibronic ground state manifold $\left|X^{1}\Sigma^{+},v=0\right\rangle$ as a function of an external electric field of magnitude $\mathcal{E}$. The hyperfine structure which is on the order of $h\times 100~{}\textrm{KHz}$ is neglected for simplicity. The lowest high and low-field seeking states (up to $\mathcal{E}$ $\approx 5~{}\mathcal{E}^{\star}$), $\left|J=0,m_{J}=0\right\rangle$ and $\left|J=1,m_{J}=0\right\rangle$, are shown as red and blue solid lines, respectively.
@@
Firstly, we consider the situation where an ion is approached by a single polar molecule prepared in its rovibronic ground state at a given distance $R$. In this situation, the charge-dipole interaction of each individual molecule in the ${}^{1}\Sigma^{+}$ ground state with the ion is determined by the Stark effect [75]:
$\hat{H}_{CD}=-\vec{d}(\vec{\mathcal{E}})\cdot\vec{\mathcal{E}}(R),$ (1)
with matrix elements $\left\langle J,m_{J}\right|-\vec{d}(\vec{\mathcal{E}})\cdot\vec{\mathcal{E}}(R)\left|J^{\prime},m_{J}^{\prime}\right\rangle$ for any two states $\left|J,m_{J}\right\rangle$ and $\left|J^{\prime},m_{J}^{\prime}\right\rangle$ coupled by $\mathcal{E}$ at a distance $R$ from the ion. The resulting mixing of rotational states with different parity gives rise to a finite permanent dipole moment in the ion-molecule center of mass frame.
The rovibrational spectrum at zero field is approximated by the Dunham series:
$U_{rov}=\sum_{lk}Y_{lk}(v+\frac{1}{2})^{l}J^{k}(J+1)^{k},$ (2)
where $Y_{lk}$ denote the Dunham coefficients and the indexes $l,k$ the powers of the vibrational and rotational quantum numbers $v$ and $J$ [76]. For the vibrational ground state $v=0$, the rotational energy is approximately $U_{rov}\approx B_{0}~{}J(J+1)$, with the rotational constant $B_{0}$. For large ion-molecule separations corresponding to the low-field limit, second order perturbation theory reveals that the shifts of the eigenvalues $U_{CD}(\mathcal{E})$ of $\hat{H}_{CD}$ are quadratic in $\mathcal{E}$. In this regime, expanding $U_{CD}$ to second order in $\mathcal{E}$, we obtain:
$U_{CD}(\mathcal{E})=U(\mathcal{E})-\left.{d(\mathcal{E})}\right|_{\mathcal{E}=0}\mathcal{E}-\frac{1}{2}\left.{\alpha(\mathcal{E})}\right|_{\mathcal{E}=0}\mathcal{E}^{2}+O(\mathcal{E}^{3}),$ (3)
@@
$d(\mathcal{E})=-\frac{\partial U_{CD}(\mathcal{E})}{\partial\mathcal{E}},~{}\alpha(\mathcal{E})=-\frac{\partial^{2}U_{CD}(\mathcal{E})}{\partial\mathcal{E}^{2}}.$
As an example, these quantities are shown for the rovibronic ground state $\left|X^{1}\Sigma^{+},v=0,J=0,m_{J}=0\right\rangle$ of NaK in Fig.4, for the corresponding eigenvalue from Fig.3. The characteristic electric field $\mathcal{E}^{\star}=B_{0}/d_{0}$ marks the transition from the weak-field to the strong-field regime. For example, in NaK this occurs at $\mathcal{E}^{\star}=2.04~{}\textrm{kV cm}^{-1}$.
Figure 4: Induced dipole moment $d(\mathcal{E})$ (blue solid line) and the static polarizability $\alpha(\mathcal{E})$ (red dashed line) for NaK as a function of the electric field $\mathcal{E}$ created by an ion separated by the distance $R$ from the (stationary) molecule. (a) Calculated $d(\mathcal{E})$ in units of the permanent dipole moment $d_{0}$ and $\alpha(\mathcal{E})$ in units of $\alpha_{0}$ (taken fom [72]) for the rovibronic ground state $\left|X^{1}\Sigma^{+},v=0,J=0,m_{J}=0\right\rangle$ and (b) the $\left|X^{1}\Sigma^{+},v=0,J=1,m_{J}=0\right\rangle$ state. The characteristic electric field $\mathcal{E}^{\star}=B_{0}/d_{0}$ is shown as the horizontal dashed line.
For $\mathcal{E}\rightarrow 0$, the linear curvature of $U_{CD}$ is negligible and $d$ vanishes, whereas $\alpha$ is maximal and the energy shifts are dominated by the quadratic Stark effect. Considering this weak-field limit (valid for $\mathcal{E}<\mathcal{E}^{\star}$) and utilizing the ion with an electric charge $q$ as a point-like source of an external electric field $\mathcal{E}=q/(4\pi\epsilon_{0}R^{2})$ for the surrounding molecules, we obtain the long-range ion-molecule potential
$U_{CD}(R)|_{R\rightarrow\infty}=-\frac{1}{2}\alpha_{0}\mathcal{E}^{2}=-\frac{1}{2}\frac{\alpha_{0}q^{2}}{(4\pi\epsilon_{0})^{2}}\frac{1}{R^{4}}.$ (4)
where the vacuum permittivity is denoted by $\epsilon_{0}$. This can be interpreted in the following way. While polar molecules possess a permanent electric dipole moment (PDM) in their own reference frame, the effective dipole moment in the laboratory frame is zero in absence of an external electric field. Therefore, in the long-range, a molecule exposed to the ion’s electric field will acquire an induced dipole moment, giving rise to an attractive potential described by Eqn.4. This result resembles the familiar expression for charge-induced dipole interactions which dominate ion-atom collisions where the long-range induction coefficient is given by $C_{4}=\alpha_{0}q^{2}/(4\pi\epsilon_{0})$ [77], and allows us to directly compare the expected long-range behavior of ions immersed into ensembles of polar molecules with the well-explored ion-atom systems. Analogously, we use the following expressions derived from the semiclassical Langevin scattering model [78, 79], to calculate the range of this potential and the characteristic energy scale for s-wave collisions:

### v2

@@ Abstract
We propose a platform for observing and controlling the interactions between atomic ions and a quantum gas of polar molecules in the ultracold regime. This approach is based on the combination of several recently developed methods in two so-far complementary research domains: ion-atom collisions and studies of ultracold polar molecules. In contrast to collisions between ions and ground-state atoms, which are dominated by losses due to three-body recombination (TBR) already at densities far below those typical for quantum degenerate ensembles, our proposal makes use of polar molecules, their rich level structure, and sensitivity to electric fields to design effective interaction potentials where ion-neutral TBR losses and molecule-molecule losses due to sticky collisions could be strongly suppressed. This may open a broad range of applications including precise control of collisional properties in molecular ensembles using ions, quantum simulations, and cold quantum chemistry between polyatomic molecules.
ion-atom interactions, molecular quantum gases, quantum chemistry
preprint: APS/123-QED
I Introduction
@@
Controlling collisions between ions and neutral atoms in a regime where quantum effects are dominant is a long-standing goal in the field of ion-atom interactions [1, 2]. This comparatively new and interdisciplinary research domain holds great promise in advancing or enabling a broad spectrum of applications in areas including quantum many-body and polaron physics [3, 4, 5, 6, 7, 8], charge mobility [9, 10], quantum simulations (QS) [11, 12], quantum chemistry (QC) [13, 14, 15, 16, 17, 18, 19], and quantum information processing (QIP) [20, 21, 22]. Most of these applications either strongly benefit from or require ultralow temperatures where collisional processes are dominated by the contribution of the lowest partial wave, the so-called s-wave regime. During the last years, enormous progress has been made towards this goal. In particular, recent advances have enabled several key achievements such as the elimination of micromotion-induced heating in generic compound optical traps [23], sympathetic cooling below the Doppler limit and close to the s-wave regime [24, 19, 23], and the controlled association of ion-atom Feshbach molecules [19].
@@
However, some substantial challenges remain an obstacle in view of many envisioned applications which require access to the deep quantum regime characterized by a combination of quantum degenerate atomic gases and ions in or close to the motional ground state. This is largely due to inelastic or reactive collisions which dominate already at moderate neutral densities [25, 23, 26] even for spin-polarized fermionic ensembles [24, 19] and lead to loss of the ion. They originate from the comparatively large characteristic range of the interaction potential of typically $R^{\star}\sim 100~{}\textrm{nm}$ between an electric charge and the induced dipole moment of the atoms exposed to the electric field of the ion. For ground-state atoms located at the distance $R$ from the ion, this attractive interaction potential scales as $1/R^{4}$ and manifests itself in the form of strongly enhanced ion-atom-atom three-body recombination processes (TBR) [25, 27, 28, 29]. At the same time, the attractive interaction is also responsible for a displacement of the ion from the node of the radiofrequency (rf) quadrupole field used to generate the confining potential for the ion, facilitating energy transfer from the rf field to the center of mass motion of the ion-atom system whenever they are allowed to undergo collisions [30].
@@
This so-called micromotion-induced heating has been identified as one of the principal obstacles to reaching the s-wave regime in conventional hybrid traps utilizing rf fields to confine the ions [30]. In order to circumvent these limitations, several strategies have been brought forward making use of different systems and mechanisms such as Rydberg atom-atom interactions [31, 32], Rydberg dressing [33], restricting the studies to few specific combinations with an extremely low reduced mass [30, 24, 19], and bichromatic traps for confining ions and atoms [23, 34]. While all these strategies have enabled recent breakthroughs, the combination of ions with quantum degenerate atomic ensembles which are characterized by large phase-space densities on the order of 1 and high densities of typically $n\sim 10^{15}~{}\textrm{cm}^{-3}$, is still out of reach. This is a consequence of the strong enhancement of TBR in vicinity of an ion in conjunction with the scaling of the related loss rates as $n^{2}$ resulting in drastically reduced ion lifetimes expected for the required densities [25, 23, 19].
@@
In parallel to the rapid evolution of studies focusing on ion-atom combinations, the complementary field of molecular quantum gases has seen a similarly impressive development. While the focus of the field shifted from chemically reactive bialkali mixtures investigated in first ground-breaking experiments [35, 36] to nonreactive combinations and directly laser-cooled molecular species, the surprisingly limited lifetimes observed for all species emerged as one of the main challenges [37, 38, 39, 40, 41, 42, 43]. For non-reactive combinations, the underlying losses have prevented the implementation of evaporative cooling, and until recently, hampered the creation of quantum degenerate gases. The exact mechanism behind these losses remains unclear and is a subject of ongoing debates [38, 39], but it is likely related to so-called sticky collisions, which occur at short intermolecular distances and involve the formation of metastable collisional complexes [37, 44, 38, 39]. With the development of methods to prevent the molecules from reaching the short range [45, 46, 47, 48, 49, 50, 51], this obstacle has recently been overcome [52, 53]. Such shielding methods rely on engineered potential barriers created through interactions with microwave [47, 48, 53], optical [51, 50] or static electric fields [54, 52]. They have enabled efficient evaporative cooling [55, 56, 57], and the first observations of molecular Fermi gases [58, 55] and Bose-Einstein Condensates (BECs) [57] have now become a reality.
@@
Here, we discuss the possibility to combine the methodology and advantageous properties of the respective systems employed in both fields in order to extend experimental studies of ion-neutral interactions at ultralow temperatures towards quantum gases of polar molecules. This platform would provide additional degrees of freedom enabling minute control over several simultaneously active interactions. We show that by exploiting these properties, characteristic quantities such as $R^{\star}$ and rates for reactive collisions could be tuned by orders of magnitude, allowing for a broad spectrum of new applications. Our article is structured in the following way: in Section II, we derive characteristic quantities of collisions between ions and polar molecules considering the influence of internal rotational degrees of freedom. We identify combinations and specific rotational states suitable for studies at ultralow collision energies. Subsequently, we discuss how existing methodology can be used to realize experimental platforms capable of combining ions and polar molecules, exemplified for Ba+-NaK, in Section III. In the rest of the paper, we consider two regimes of neutral molecule density, that is, dilute molecular gases suitable for novel studies of quantum chemistry (Section IV), and dense quantum gases where the interplay of attractive and repulsive long-range and short-range interactions is expected to play an important role (Section V). Lastly, we discuss potential applications of the proposed hybrid ion-molecule system in studies of quantum-chemical reactions, many-body phenomena, and dynamical processes in molecular quantum gases as well as their control in Section VI.
II Polar molecules in the electric field of an atomic ion
II.1 Molecules in the rovibronic ground state
@@
Figure 1: (top panel) Magnitude of the electric field emanating from a single ion as a function of its distance to an approaching neutral molecule. The grey dashed line denotes the characteristic electric field of NaK, $\mathcal{E}^{\star}$(NaK)$=B_{0}/d_{0}=2.04~{}\textrm{kV cm}^{-1}$. (bottom panel) Numerically calculated energy levels of the NaK molecule for the lowest rotational states $\left|J=0,~{}1,~{}2,~{}3;m_{J}\right\rangle$ (considering a maximal $J_{max}=10$) within the vibronic ground state manifold $\left|X^{1}\Sigma^{+},v=0\right\rangle$ as a function of an external electric field of magnitude $\mathcal{E}$. The hyperfine structure which is on the order of $h\times 100~{}\textrm{kHz}$ is neglected for simplicity. The lowest high and low-field seeking states (up to $\mathcal{E}$ $\approx 5~{}\mathcal{E}^{\star}$), $\left|J=0,m_{J}=0\right\rangle$ and $\left|J=1,m_{J}=0\right\rangle$, are shown as red and blue solid lines, respectively.
@@
Firstly, we consider the situation where an ion is approached by a single polar molecule prepared in its rovibronic ground state $\left|X^{1}\Sigma^{+},v=0,J=0,m_{J}=0\right\rangle$ at a given distance $R$, with $v$, $J$ and $m_{J}$ denoting the vibrational and rotational quantum number and the projection of $\vec{J}$ onto the electric field $\vec{\mathcal{E}}$, respectively. This situation can be treated in the same way as the interaction of induced dipoles in a radially symmetric external electric field created by a point-like charge. The charge-dipole interaction of each individual molecule with the ion is determined by the Stark effect [59]:
$\hat{H}_{CD}=-\vec{d}(\vec{\mathcal{E}})\cdot\vec{\mathcal{E}}(R).$ (1)
The competition between the molecule’s rotation and the alignment of its dipole moment $\vec{d}(\vec{\mathcal{E}})$ can be expressed in terms of the characteristic electric field $\mathcal{E}^{\star}=B_{0}/d_{0}$, where $B_{0}$ is the rotational constant and $d_{0}$ the permanent dipole moment (PDM) of the molecule in the body-fixed frame. The rotational energy in the vibrational ground state $v=0$ is approximately $U_{rot}\approx B_{0}~{}J(J+1)$. For large ion-molecule separations corresponding to the low-field limit where $\mathcal{E}\ll\mathcal{E}^{\star}$ the eigenvalues $U_{CD}(\mathcal{E})$ of $\hat{H}_{CD}$ are quadratic in $\mathcal{E}$:
$U_{CD}(\mathcal{E})=U_{rot}-\left.{d(\mathcal{E})}\right|_{\mathcal{E}=0}\mathcal{E}-\frac{1}{2}\left.{\alpha(\mathcal{E})}\right|_{\mathcal{E}=0}\mathcal{E}^{2}+O(\mathcal{E}^{3}),$ (2)
@@
$d(\mathcal{E})=-\frac{\partial U_{CD}(\mathcal{E})}{\partial\mathcal{E}},~{}\alpha(\mathcal{E})=-\frac{\partial^{2}U_{CD}(\mathcal{E})}{\partial\mathcal{E}^{2}}.$ (3)
As an example, these quantities are shown for the rovibronic ground state of NaK in Fig.2, for the corresponding eigenvalues from Fig.1.
Figure 2: Induced dipole moment $d(\mathcal{E})$ (blue solid line) and the static polarizability $\alpha(\mathcal{E})$ (red dashed line) for NaK as a function of the electric field $\mathcal{E}$ created by an ion separated by the distance $R$ from the (stationary) molecule. (a) Calculated $d(\mathcal{E})$ in units of the permanent dipole moment $d_{0}$ and $\alpha(\mathcal{E})$ in units of $\alpha_{0}$ (taken fom [60]) for the rovibronic ground state $\left|X^{1}\Sigma^{+},v=0,J=0,m_{J}=0\right\rangle$ and (b) the $\left|X^{1}\Sigma^{+},v=0,J=1,m_{J}=0\right\rangle$ state. The characteristic electric field $\mathcal{E}^{\star}$ is shown as the horizontal dashed line.
Utilizing the ion with an electric charge $q$ as a point-like source of an external electric field $\mathcal{E}=q/(4\pi\epsilon_{0}R^{2})$ for the surrounding molecules, we obtain the long-range ion-molecule potential
$U_{CD}(R)|_{R\rightarrow\infty}=-\frac{1}{2}\alpha_{0}\mathcal{E}^{2}=-\frac{1}{2}\frac{\alpha_{0}q^{2}}{(4\pi\epsilon_{0})^{2}}\frac{1}{R^{4}},$ (4)
whith the vacuum permittivity $\epsilon_{0}$. This result is identical with the familiar expression for charge-induced dipole interactions which dominate ion-atom collisions where the long-range induction coefficient is given by $C_{4}=\frac{1}{2}\alpha_{0}q^{2}/(4\pi\epsilon_{0})^{2}$ [61]. This allows us to calculate $C_{4}$ coefficients for polar molecules and to directly compare the expected long-range behavior with the well-explored ion-atom systems. Analogously, we use the following expressions derived from the semiclassical Langevin scattering model [62, 63], to calculate the range of this potential and the characteristic energy scale for s-wave collisions: